OpenStudy (anonymous):

How do I find the adjoint of this matrix? (1 3 1) (2 1 1) (-2 2 -1)

OpenStudy (amistre64):

cofactor transpose = adjoint

OpenStudy (amistre64):

\[\begin{array}c +\left|\begin{matrix}1&1\\2&-1 \end{matrix} \right|& -\left|\begin{matrix}2&1\\-2&-1 \end{matrix} \right|& +\left|\begin{matrix}2&1\\-2&2 \end{matrix} \right| \\\\ -\left|\begin{matrix}3&1\\2&-1 \end{matrix} \right|& +\left|\begin{matrix}1&1\\-2&-1 \end{matrix} \right|& -\left|\begin{matrix}1&3\\-2&2 \end{matrix} \right| \\\\ +\left|\begin{matrix}3&1\\1&1 \end{matrix} \right|& -\left|\begin{matrix}1&1\\2&1 \end{matrix} \right|& +\left|\begin{matrix}1&3\\2&1 \end{matrix} \right| \\\\ \end{array}\] i just cant recall if we include the "entry" like in a determinant at the moment

OpenStudy (amistre64):

nah, thats right

OpenStudy (amistre64):

all those are little determinants; number them out and transpose this thing :)

OpenStudy (amistre64):

-3 0 6 5 1 -8 ; chk my math to be sure, but then we transpose 2 1 -5 -3 5 2 0 1 1 ; would be the adjoint if I did it right :) 6 -8 -5